Gigabyte, Inc. manufactures three products for the computer industry: Gismos (product G): annual sales, 8,000 units ThingamajDirect-labor budget (based on budgeted annual sales): Product G: Product T: Product W: Total 8,000 X $ 16.00 15,000 x $12.00Required: 1. Is product G the companys least profitable product? 2. Is product W a profitable product for Gigabyte, Inc.? 4.Complete this question by entering your answers in the tabs below. Required 1 Required 2 Required 4 Required 5 Required 6 ShoComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 4 L Required 5 Required 6 DComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 4 Required 5 Required 6 Bas
Answer #1

Solution:

4. Determination of Percentages for Raw-Material costs
Product Raw Material cost per unit Annual Volume Annual Raw Material Cost % of Total Raw Material Cost
G $ 35.00 8,000 280,000 25%
T $ 52.50 15,000 787,500 69%
W $ 17.50 4,000 70,000 6%
Total 1,137,500 100%
5. Computation of Product Cost on Activity Based Costing System
Particulars Product G Product T Product W
Direct Material 280,000.00 787,500.00 70,000.00
Direct Labour 128,000.00 180,000.00 32,000.00
Machine Setup 1,050.00 1,575.00 2,625.00
Machinery 306,250.00 612,500.00 306,250.00
Inspection 78,750.00 236,250.00 210,000.00
Material handling 218,750.00 603,750.00 52,500.00
Engineering 120,662.50 34,475.00 189,612.50
Total Product Cost 1,133,462.50 2,456,050.00 862,987.50
6. Calculation of Target Price for each product
Particulars Product G Product T Product W
Target Price based on new product cost $ 212.52 $ 245.61 $ 323.63

Workings:

Target Price based on new product cost = (Total Product Cost / Annual Volume)* 150%
Product G = (1,133,462.50 / 8,000) * 150%
= $ 212.52
Product T = (2,456,050 / 15,000) * 150%
= $ 245.61
Product W = (862,987.50 / 4,000) * 150%
= $ 323.63